OpenStudy (anonymous):

Can someone please walk me through how to do square roots like breaking it down into a a number and a radical? Can you help me by giving me a few sample questions and walking me through it please? thanks!

OpenStudy (lgbasallote):

do you mean \[\huge \sqrt{48} \implies \sqrt{16 \times 3} \implies 4\sqrt 3\] something like that? or no?

OpenStudy (anonymous):

yes something like that! please help

OpenStudy (lgbasallote):

i think one of these two people typing something will give something helpful..

OpenStudy (unklerhaukus):

factorise the number in to a product of primes \[48=2\times 24=2^2\times12=2^3\times6=2^4\times3=(2^2)^2\times3\] \[\sqrt{48}=\sqrt{(2^2)^2\times3}=2^2\sqrt3=4\sqrt3\]

OpenStudy (lgbasallote):

@sparky16 try checking this link out then tell me if it helps or not openstudy.com/updates/4fbae0ebe4b05565342fe421

OpenStudy (lgbasallote):

http://openstudy.com/updates/4fbae0ebe4b05565342fe421

OpenStudy (lgbasallote):

^there

OpenStudy (valpey):

So if what @lgbasallote and @UnkleRhaukus wrote isn't exactly what you were getting at the missing step is:\[\large\sqrt{48}\Rightarrow \sqrt{16*3}\Rightarrow\sqrt{4^2*3}\Rightarrow\sqrt{4^2}*\sqrt{3}\Rightarrow 4*\sqrt{3}\]

OpenStudy (lgbasallote):

hehe i was just asking about the question...i wasnt going in any details but good eye in seeing the missing step :D

OpenStudy (anonymous):

ok to clarify.. u pull out a group of two numbers and if there are four number you times them together to put on the outside? and what if there are 3 groups of two numbers... i dont know if this is possible but like 2 groups of 2's and one group of 3's

OpenStudy (unklerhaukus):

lets see if you can do this one\[\sqrt{8100}=\]

OpenStudy (unklerhaukus):

step 1 would be to factorise 8100

OpenStudy (valpey):

And it will always be true for any a and b that: \[\sqrt{a*a*b}=\sqrt{a^2b}=\sqrt{a^2}\sqrt{b}=a\sqrt{b}\]And sometimes b itself will have more ways to factor it and you can work to simplify \[\sqrt{b}\]

OpenStudy (unklerhaukus):

tue for real factors yes

OpenStudy (lgbasallote):

@sparky16 do you mean something like \[\sqrt{2 \times 2 \times 3 \times 3}\] if so pull out 2 and 3 \[\implies 2(3) \sqrt{\cancel{2 \times 2 \times 3 \times 3}} \]

OpenStudy (lgbasallote):

\[\sqrt{2 \times 2 \times 2 \times 2 \times 3 \times 3} \] \[\implies 2\sqrt{\cancel{2 \times 2} \times 2 \times 2 \times 3 \times 3}\] \[\implies 2(2) \sqrt{ \cancel{2 \times 2 \times 2 \times 2 \times 2} \times 3 \times 3}\] \[\implies 2(2)(3)\sqrt{\cancel{2 \times 2\times 2\times 2\times 3 \times 3}}\] does that help?

OpenStudy (anonymous):

@igbasallote- like 2 x 2 x 2 x 2 x 3 x 3 ( just times them all together right?) or do you take about the group of 2 and just time it by 3 to get 6 or is 12 because you times all of them?

OpenStudy (lgbasallote):

uhh i think i exceeded by one 2 in the square root in the second line...iignore that typo

OpenStudy (lgbasallote):

it's 12...you multiply everything you pulled out

OpenStudy (anonymous):

okay thankyou and unklerhaukus.... when there are multiple pairs to pull out what do i do?

OpenStudy (anonymous):

8100 has 5 pairs to pull out so what stays inside the radical?

OpenStudy (unklerhaukus):

\[\sqrt{8100}=\sqrt{81\times100}\]\[=\sqrt{9^2\times10^2}=\sqrt{9^2}\times\sqrt{10^2}\]\[=9\times10=90\]

OpenStudy (anonymous):

i did the factor tree and i got 90 just like you thank you... and one last question... when dealing with the number inside the radical, do i just times together the leftover numbers that aren't pairs?

OpenStudy (lgbasallote):

yes @sparky16 same thing i did in that tutorial i showed you

OpenStudy (unklerhaukus):

or you could have done it like this \[\sqrt{8100}=\sqrt{2^2\times3^4\times5^2}=\sqrt{2^2\times(3^2)^2\times5^2}\] \[=\sqrt{2^2}\times\sqrt{(3^2)^2}\times\sqrt{5^2}\]\[=2\times3^2\times5\]

OpenStudy (anonymous):

thank you so much to everyone that helped! I understand it now! Next time I log on I may have another question in the future so look for my name cuz u were all amazing help!(:

OpenStudy (lgbasallote):

haha tag us :p lol

OpenStudy (anonymous):

alright will do when I need help! especially you igbasalote! thanks! kk bye